OpenStudy (anonymous):

Please help! If sin Θ = 3 over 7 and tan Θ < 0, what is the value of cos Θ?

OpenStudy (anonymous):

Since \[\sin \Theta = \frac{ opposite }{ hypotenuse }\] \[\cos \Theta = \frac{ adjacent }{ hypotenuse }\] \[\tan \Theta = \frac{ opposite }{ adjacent }\] you can draw the triangle:  Then you can use Pythagorean theorem to find the adjacent side and therefore the cosine